x -2 -1 0 1 2
y .25 .5 1 2 4
Given the table above, graph the function, identify the graph of the function, explain your choice, and give the domain and range as shown in the graph, and also the domain and range of the entire function.
Help I am so confused with this!!!
the table doesn't show up but its 6 columns and 2 rows.

Answer by stanbon(48535)

x -2 -1 0 1 2
y .25 .5 1 2 4
Given the table above, graph the function, identify the graph of the function, explain your choice, and give the domain and range as shown in the graph, and also the domain and range of the entire function.
Notice that as the x-values go up by one the
y-values go up as multiples of two.
So the equation is an exponential function
of the form y=ab^x.
You need to find a and b.
-------------------------------
Using the points (0,1) and (2,4)
you get
1=ab^0
1=a(1)
a=1
-----------
4=(1)b^2
b=2
----------
So the equation you want is y=2^x
The domain of the function is all real numbers.
The range is all real numbers greater than zero.
Cheers,
Stan H.